# Step-by-Step Mastery: Enhancing Soft Constraint Following Ability of Large Language Models

### v3

@@ Abstract
It is crucial for large language models (LLMs) to follow instructions that involve multiple constraints. However, it is an unexplored area to enhance LLMs’ ability to follow soft constraints. To bridge the gap, we initially design a pipeline to construct datasets with high-quality outputs automatically. Additionally, to fully utilize the positive and negative samples generated during the data construction process, we choose Direct Preference Optimization (DPO) as the training method. Furthermore, taking into account the difficulty of soft constraints indicated by the number of constraints, we design a curriculum learning training paradigm based on the constraint quantity. We experimentally evaluate the effectiveness of our methods in improving LLMs’ soft constraint following ability and analyze the factors driving the improvements. The datasets and code are publicly available at https://github.com/Rainier-rq/FollowSoftConstraint.
@@
Qingyu Ren1,3, Jie Zeng1, Qianyu He1, Jiaqing Liang2†, Yanghua Xiao1† Weikang Zhou3, Zeye Sun3, Fei Yu3 1Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University 2School of Data Science, Fudan University 3Ant Group {qyren24, jzeng23, qyhe21}@m.fudan.edu.cn, {liangjiaqing, shawyh}@fudan.edu.cn
@@ 1 Introduction
Following instructions with soft constraints is imperative for LLMs Jiang et al. (2023b); Qin et al. (2024). Constraints can be categorized into soft and hard constraints. Hard constraints can be explicitly expressed as specific rules and directly verified through programming methods Zhou et al. (2023a); He et al. (2024a). For example, Python can parse JSON data to verify whether it follows specific format constraints. However, instructions in real-world applications often contain semantic-level limitations, which can be categorized as soft constraints. Soft constraints include restrictions related to content Liang et al. (2024); Zhang et al. (2023), specific backgrounds Shanahan et al. (2023); Liu et al. (2023), and the style of expression Sigurgeirsson and King (2024); Mukherjee et al. (2024). They are difficult to verify automatically through programming methods. As shown in Fig. 1, following soft constraints is challenging for LLMs.
@@
Figure 2: The framework of our study. We first design a pipeline to automatically construct datasets with high-quality outputs for soft constraint following. Then, we propose a DPO training paradigm based on curriculum learning. CL denotes curriculum learning.
@@
Improving LLMs’ soft constraint following ability is under-explored. Much of the existing work focuses on hard constraints, and it is challenging for LLMs to generalize from hard to soft constraints He et al. (2024a). Additionally, existing research focuses on evaluating LLMs’ soft constraint following ability Chen et al. (2024a); Qin et al. (2024) rather than improving the ability. Overall, existing work on soft constraints has following three limitations: First, directly using advanced models to generate outputs for constructing the dataset results in low quality Xu et al. (2023); Chiang et al. (2023). Soft constraints are especially challenging for these models. On FollowBench Jiang et al. (2023b), GPT-4 demonstrates a hard satisfaction rate of merely 74.4%, making the assurance of high-quality outputs difficult. Second, simply using Supervised Fine-Tuning (SFT) Ouyang et al. (2022) to train the model leads to poor efficiency and performance. Slight changes about the soft constraint in the instruction can result in significant variations in the output He et al. (2024a). It is hard to learn these differences just from positive samples. Third, existing work ignores the difficulty of soft constraints indicated by the number of constraints. Many studies show that organizing training data in the order of the curriculum is better than random shuffling Sun et al. (2024); Lee et al. (2024). Therefore, an efficient training paradigm is required.
@@
In this work, we systematically investigate strategies to enhance the ability of LLMs to follow soft constraints, with the framework shown in Fig. 2, including constructing datasets with high-quality outputs and proposing a DPO training paradigm based on curriculum learning. To construct datasets with high-quality outputs, we progressively add constraints to the instructions and incorporate Judger to reorder the outputs based on constraint following. To fully leverage the positive and negative samples generated during Judger reordering, we apply Direct Preference Optimization (DPO) Rafailov et al. (2024) as the training method. By contrasting positive and negative outputs, it can capture how minor soft constraint adjustments affect output preferences. To account for the difficulty of soft constraints indicated by the number of constraints, we propose a training paradigm that constructs curricula based on the number of constraints in the instruction. In this paradigm, the model starts by learning simpler tasks (fewer constraints) and progresses to more complex ones (more constraints). Our methods can improve LLMs’ soft constraint following ability effectively.
@@
Our contributions are summarized as follows: (1) We design a pipeline to automatically construct datasets with high-quality outputs for soft constraint following. (2) We introduce a DPO training paradigm that constructs curricula based on the number of constraints. (3) We conduct extensive experiments to validate the effectiveness of our methods and analyze the reasons for the performance improvements.
@@ 2 Related Work
Soft Constraint Following Existing research on soft constraint following focuses on evaluating the ability of LLMs by constructing benchmarks Jiang et al. (2023b); Qin et al. (2024). These benchmarks include a variety of fine-grained constraint types Zhang et al. (2024). These constraints can be categorized into several types: (1) Content soft constraints involve restrictions on the scope or depth of the responses Zhou et al. (2023b); Ashok and Poczos (2024). (2) Situation soft constraints refer to the background limitations Wang et al. (2023b); Shao et al. (2023). (3) Style soft constraints limit the manner of expressions Tao et al. (2024); Pu et al. (2024). Some works directly utilize responses generated by GPT-4 to construct datasets Sun et al. (2024); Peng et al. (2023). However, LLMs struggle to follow soft constraints He et al. (2024b), making responses unreliable. Different from these, we construct datasets with high-quality outputs to improve LLMs’ soft constraint following ability.
@@ 3.1 Dataset Construction
We first construct a multi-constraint instruction following dataset. Adding all constraints at once increases the complexity of the instructions rapidly, making it difficult for the model to understand all constraints He et al. (2024a). To addresss this, we propose a progressive construction method, adding one constraint at a time, which allows the model to understand and learn to follow each constraint effectively. Moreover, existing works in dataset construction rely on advanced models to directly generate the outputs Sun et al. (2024). However, even GPT-4 is struggling to follow the instructions with complex soft constraints Jiang et al. (2023b); Qin et al. (2024). To obtain high-quality outputs, we use Judger to reorder the outputs based on the extent of constraint following. Overall, our pipeline consists of two successive steps: Progressive Construction and Judger Reordering.
@@ 3.3.1 Data Statistics
Method Nums. Cons. Pair. Prog.
@@
Suri Pham et al. (2024) 20000 S $\checkmark$ $\times$
AutoIF Dong et al. (2024) - H $\checkmark$ $\times$
Complex to Simple He et al. (2024a) 12939 H $\checkmark$ $\times$
@@
Table 2: A detailed comparison of related works. Ours represents our dataset. “Nums.”, “Cons.”, “Pair.”, and “Prog.”  denote the number of preference pairs, constraint types, whether to perform pairwise comparison, and whether the dataset is progressively constructed.
@@ 4 Experiments
Model BaseModel FollowBench (HSR) IFEval
@@
Qwen2-72B-Instruct Yang et al. (2024) Qwen 82.4 68.2 62.6 61.2 54.7 65.8 77.1 84.4 80.4 86.9 82.2
@@
Vicuna-13B-v1.5 Chiang et al. (2023) LLaMA2 56.2 42.9 32.3 32.1 24.6 37.6 43.1 53.6 46.6 58.0 50.3
Mistral-7B-ShareGPT Jiang et al. (2023a) Mistral 51.6 45.8 38.3 25.8 20.7 36.5 43.6 53.5 47.3 57.8 50.6
LLaMA3-8B-ShareGPT Dubey et al. (2024) LLaMA3 57.3 49.0 42.3 30.5 27.8 41.4 37.7 49.5 42.3 53.8 45.8
Conifer-7B Sun et al. (2024) Mistral 54.3 49.5 49.3 40.8 30.5 44.9 45.8 57.1 50.8 62.0 53.9
@@
Mistral-7B-Instruct-v0.3SFT+Judger Mistral 58.7 52.4 42.5 37.2 35.6 45.3 56.8 67.8 60.6 71.3 64.1
Mistral-7B-Instruct-v0.3DPO+Judger+CL Mistral 63.3 56.0 47.5 40.0 36.7 48.7 53.4 63.5 57.1 67.5 60.4
@@
LLaMA2-13B-Chat-HFSFT+Judger LLaMA2 53.6 43.0 41.2 34.9 25.8 39.7 32.3 44.7 43.1 55.0 43.8
LLaMA2-13B-Chat-HFDPO+Judger+CL LLaMA2 53.6 48.3 42.5 29.2 29.9 40.7 35.1 46.6 48.2 58.9 47.2
[…]
LLaMA3-8B-InstructSFT+Judger LLaMA3 66.3 55.4 50.1 49.7 39.8 52.3 70.4 77.8 73.2 80.1 75.4
LLaMA3-8B-InstructDPO+Judger+CL LLaMA3 69.2 59.6 50.8 48.9 44.6 54.6 72.5 80.3 77.1 84.1 78.5
Table 3: The overall performance on FollowBench and IFEval. We use bold for the best results and underlined for the second-best results among the models ranging from 7B to 13B parameter sizes. indicates that the results are directly sourced from the original benchmarks.
@@ 4.1 Experiment Setup
Models. We conduct experiments on three widely recognized base LLMs, LLaMA3-8B-Instruct Dubey et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023a), and LLaMA2-13B-Chat-HF Touvron et al. (2023). Within our experimental framework, we compare three approaches: (1) BASE directly utilizes the base model to generate outputs. (2) SFT+Judger applies supervised fine-tuning on LLMs using the instruction-response pairs $(I_{n},O_{w_{n}})$ generated by Judger (§3.1.13.1.2). (3) DPO+Judger+CL utilizes Judger to obtain high-quality training data, which is then used for DPO training following the curriculum learning paradigm (§3.1.2, §3.2).
For baseline comparisons, we select various open-source and proprietary LLMs. Among the proprietary models, we include GPT-4 Achiam et al. (2023) and GPT-3.5 Turbo. Additionally, we compare our approach with several open-source LLMs, including models specifically trained to improve general instruction following abilities and models focusing on enhancing the ability to follow complex instructions. We also compare our models with two 70B-sized powerful models.
@@
IFEval Zhou et al. (2023a) is a benchmark designed to evaluate the ability to follow hard constraints. It defines 25 distinct types of verifiable instructions, each containing between 1 and 3 constraints. FollowBench Jiang et al. (2023b) is a benchmark that evaluates the ability of models to follow both soft and hard constraints across multiple levels of granularity.
@@ 4.2 Main Results
As shown in Tab. 4, our method significantly enhances the model’s ability to follow soft constraints. Compared to the BASE method, there is a significant performance improvement across both benchmarks when three models are trained using the DPO+Judger+CL method, particularly on IFEval. The improvement is significant on difficult soft constraint following tasks, especially at the L5 difficulty level in FollowBench. Specifically, LLaMA-3-8B-Instruct shows an improvement of 5.5% at the L5 difficulty level.
After training with the DPO+Judger+CL method, within the range of models with 8B parameters, Mistral-7B-Instruct-v0.3 and LLaMA-3-8B-Instruct outperform other models on both benchmarks. Within the 13B model category, LLaMA2-13B-Chat-HF outperforms WizardLM-v1.2-13B and Vicuna-13B-v1.5 on FollowBench. Although its performance on IFEval, which only includes hard constraints, is lower than that of the other two models, our method demonstrates superior results on FollowBench, a more complex task that incorporates both soft and hard constraints.
After training with the SFT+Judger method, there is a decline in the performance of LLaMA2-13B-Chat-HF and Mistral-7B-Instruct-v0.3. This drop is attributed to the model’s integration of various specialized training techniques during its initial training phase. Although the SFT+Judger method performs better than DPO+Judger+CL on IFEval for the Mistral-7B-Instruct-v0.3 model, the method even performs worse than the BASE method on FollowBench, failing to effectively improve the model’s ability to follow soft constraints. Moreover, the performance of the SFT+Judger method is worse than the DPO+Judger+CL method on both benchmarks for the other two models.

### v4

@@ Abstract
It is crucial for large language models (LLMs) to follow instructions that involve multiple constraints. In real-world scenarios, user instructions often contain soft constraints, which are semantically related and cannot be rule-based verified, posing challenges for LLMs. To enhance the soft constraint following ability of LLMs, we initially design a pipeline to construct datasets with high-quality outputs for instructions containing soft constraints automatically. Additionally, to fully utilize the positive and negative samples generated during the data construction process, we choose Direct Preference Optimization (DPO) as the training method. Furthermore, taking into account the difficulty of soft constraints indicated by the number of constraints, we design a curriculum learning training paradigm based on the constraint quantity. We experimentally evaluate the effectiveness of our methods in improving LLMs’ soft constraint following ability and analyze the factors driving the improvements. The datasets and code are publicly available at https://github.com/Rainier-rq/FollowSoftConstraint.
@@
Qingyu Ren1,3, Jie Zeng1, Qianyu He1, Jiaqing Liang2thanks:  Corresponding author., Yanghua Xiao1 Weikang Zhou3, Zeye Sun3, Fei Yu3 1Shanghai Key Laboratory of Data Science, College of Computer Science and Artificial Intelligence, Fudan University 2School of Data Science, Fudan University 3Ant Group {qyren24, jzeng23, qyhe21}@m.fudan.edu.cn, {liangjiaqing, shawyh}@fudan.edu.cn
@@ 1 Introduction
Following instructions with soft constraints is imperative for LLMs Jiang et al. (2023b); Qin et al. (2024). Constraints can be categorized into soft and hard constraints. Hard constraints can be explicitly expressed as specific rules and directly verified through programming methods Zhou et al. (2023a); He et al. (2024). For example, Python can parse JSON data to verify whether it follows specific format constraints. However, instructions in real-world applications often contain semantic-level limitations, which can be categorized as soft constraints. Soft constraints include restrictions related to content Liang et al. (2024); Zhang et al. (2023), specific backgrounds Shanahan et al. (2023); Liu et al. (2023), and the style of expressions Sigurgeirsson and King (2024); Mukherjee et al. (2024). They are difficult to verify automatically through programming methods. As shown in Fig. 1, following soft constraints is challenging for LLMs.
@@
Figure 2: The framework of our study. We first design a pipeline to automatically construct datasets with high-quality outputs for soft constraint following. Then, we propose a reinforcement learning training paradigm based on curriculum learning. CL denotes curriculum learning.
@@
Existing methods for improving soft constraint following ability of LLMs have following limitations: First, much of the existing work focuses on evaluating LLMs’ soft constraint following ability Chen et al. (2024a); Qin et al. (2024) rather than how to improve the ability. Second, when constructing soft constraint datasets, current methods typically introduce all constraints at once and directly prompt advanced models to generate responses Xu et al. (2023); Chiang et al. (2023). However, these responses may not satisfy all constraints. For example, GPT-4 demonstrates a constraint satisfaction rate of only 74.4% on FollowBench Jiang et al. (2023b), making the assurance of high-quality outputs difficult. Third, in terms of the training paradigm, existing work ignores the difficulty of constraint following indicated by the number of constraints He et al. (2024); Qi et al. (2024). Many studies show that organizing training data in a curriculum-based order—starting from simpler examples and gradually increasing complexity—leads to better performance than random shuffling Sun et al. (2024); Lee et al. (2024). Therefore, a pipeline constructing datasets with high-quality outputs and an efficient training paradigm are required.
@@
In this work, we systematically investigate strategies to enhance the ability of LLMs to follow soft constraints, with the framework shown in Fig. 2, including constructing datasets with high-quality outputs and proposing a reinforcement learning training paradigm based on curriculum learning. To construct datasets with high-quality outputs for soft constraint following, we progressively add constraints to the instructions and incorporate Judger to reorder the outputs based on constraint following. To fully leverage the positive and negative samples generated during Judger reordering, we apply Direct Preference Optimization (DPO) Rafailov et al. (2024) as the training method. Since soft constraints cannot be verified through rule-based methods, Judger reordering can ensure that the constructed preference dataset adheres to the correct constraint-following behavior. To account for the difficulty of soft constraints indicated by the number of constraints, we propose a training paradigm that constructs curricula based on the number of constraints in the instruction. In this paradigm, the model starts by learning simpler tasks (fewer constraints) and progresses to more complex ones (more constraints). Our methods can improve LLMs’ soft constraint following ability effectively. Moreover, our method is also effective for hard constraints.
@@
Our contributions are summarized as follows: (1) We design a pipeline to automatically construct datasets with high-quality outputs for soft constraint following. (2) We introduce a reinforcement learning training paradigm that constructs curricula based on the number of constraints. (3) We conduct extensive experiments to validate the effectiveness of our methods and analyze the reasons for the performance improvements.
@@ 2 Related Work
Soft Constraint Following Existing research on soft constraint following focuses on evaluating the ability of LLMs by constructing benchmarks Jiang et al. (2023b); Qin et al. (2024). These benchmarks include a variety of fine-grained constraint types Zhang et al. (2024). These constraints can be categorized into several types: (1) Content soft constraints involve restrictions on the scope or depth of the responses Zhou et al. (2023b); Ashok and Poczos (2024). (2) Situation soft constraints refer to the background limitations Wang et al. (2023b); Shao et al. (2023). (3) Style soft constraints limit the manner of expressions Tao et al. (2024); Pu et al. (2024). Some works directly utilize responses generated by GPT-4 to construct datasets Sun et al. (2024); Peng et al. (2023). Different from these, we propose a pipline constructing datasets with high-quality outputs to improve LLMs’ soft constraint following ability.
@@ 3.1 Dataset Construction
We first construct a multi-constraint instruction following dataset. Adding all constraints at once increases the complexity of the instructions rapidly, making it difficult for the model to understand all constraints He et al. (2024). To addresss this, we propose a progressive construction method, adding one constraint at a time, which allows the model to understand and learn to follow each constraint effectively. Moreover, existing works in dataset construction rely on advanced models to directly generate the outputs Sun et al. (2024). However, even GPT-4 is struggling to follow the instructions with complex soft constraints Jiang et al. (2023b); Qin et al. (2024). Since soft constraints cannot be verified for compliance through rules, we use Judger to reorder the outputs based on the extent of constraint following to obtain high-quality outputs. Overall, our pipeline consists of two successive steps: Progressive Construction and Judger Reordering.
@@ 3.3.1 Data Statistics
Method Nums. Cons. Reor. Prog.
@@
Suri Pham et al. (2024) 20000 S $\times$ $\times$
AutoIF Dong et al. (2024) - H $\times$ $\times$
Complex to Simple He et al. (2024) 12939 H $\times$ $\times$
@@
Table 2: A detailed comparison of related works. Ours represents our dataset. “Nums.”, “Cons.”, “Reor.”, and “Prog.”  denote the number of preference pairs, constraint types, whether to perform output reordering, and whether the dataset is progressively constructed.
@@ 4 Experiments
Model FollowBench (HSR) IFEval
@@
Vicuna-13B-v1.5 Chiang et al. (2023) 56.2 42.9 32.3 32.1 24.6 37.6 43.1 53.6 46.6 58.0 50.3
Conifer-7B Sun et al. (2024) 54.3 49.5 49.3 40.8 30.5 44.9 45.8 57.1 50.8 62.0 53.9
Conifer-7B-DPO Sun et al. (2024) 60.3 53.6 48.0 47.1 41.0 50.0 48.1 59.1 52.3 63.3 55.7
MistralCRAB+DPO Qi et al. (2024) 66.1 53.6 53.4 42.4 31.7 49.4 49.7 61.5 57.7 68.5 59.3
Mistral-7B-ShareGPT Jiang et al. (2023a) 51.6 45.8 38.3 25.8 20.7 36.5 43.6 53.5 47.3 57.8 50.6
Llama3CRAB+DPO Qi et al. (2024) 64.6 49.0 41.6 35.8 36.8 45.5 40.3 52.0 47.7 58.9 49.7
@@
Mistral-7B-Instruct-v0.2SFT+Judger 58.1 50.5 44.6 36.0 36.2 45.1 44.2 56.5 47.5 60.2 52.1
Mistral-7B-Instruct-v0.2DPO+Judger+CL 58.8 54.8 46.3 39.4 35.0 46.9 46.6 58.5 52.1 63.7 55.2
@@
Mistral-7B-Instruct-v0.3SFT+Judger 58.7 52.4 42.5 37.2 35.6 45.3 56.8 67.8 60.6 71.3 64.1
Mistral-7B-Instruct-v0.3DPO+Judger+CL 63.3 56.0 47.5 40.0 36.7 48.7 53.4 63.5 57.1 67.5 60.4
@@
LLaMA3-8B-InstructSFT+Judger 66.3 55.4 50.1 49.7 39.8 52.3 70.4 77.8 73.2 80.1 75.4
LLaMA3-8B-InstructDPO+Judger+CL 69.2 59.6 50.8 48.9 44.6 54.6 72.5 80.3 77.1 84.1 78.5
Table 3: The overall performance on FollowBench and IFEval. We use bold for the best results and underlined for the second-best results among the models ranging from 7B to 13B parameter sizes.
@@ 4.1 Experiment Setup
Models. We conduct experiments on several widely recognized base LLMs, including the LLaMA series Dubey et al. (2024), Mistral series Jiang et al. (2023a), and Qwen2.5 series Yang et al. (2024). Within our experimental framework, we compare three approaches: (1) BASE directly utilizes the base model to generate outputs. (2) SFT+Judger applies supervised fine-tuning on LLMs using the instruction-response pairs $(I_{n},O_{w_{n}})$ generated by Judger (§3.1.13.1.2). (3) DPO+Judger+CL utilizes Judger to obtain high-quality training data, which is then used for DPO training following the curriculum learning training paradigm (§3.1.2, §3.2). For baseline comparisons, we include proprietary models such as GPT-4 Achiam et al. (2023) and GPT-3.5 Turbo. Additionally, we also select models specifically designed to enhance the ability to follow multi-constraint instructions.
@@
FollowBench Jiang et al. (2023b) is a benchmark that evaluates the ability of models to follow both soft and hard constraints across multiple levels of granularity. CFBench Zhang et al. (2024) is a benchmark that evaluates the ability of models to follow soft constraints. Each example needs to be evaluated by GPT-4. IFEval Zhou et al. (2023a) is a benchmark designed to evaluate the ability to follow hard constraints.
Figure 4: Results across various constraint categories.
Model Hard Satisfaction Rate Soft Satisfaction Rate
L1 L2 L3 L4 L5 Avg L1 L2 L3 L4 L5 Avg
Mistral-7B-Instruct-v0.2BASE 70.5 58.4 50.6 51.7 39.0 54.0 70.5 67.5 66.1 64.1 61.2 65.9
Mistral-7B-Instruct-v0.2DPO+Judger+CL 70.3 65.4 55.8 47.4 41.5 56.1 70.3 71.3 67.1 64.1 63.9 67.3
[…]
LLaMA3-8B-InstructDPO+Judger+CL 81.2 67.5 64.9 56.3 49.9 64.0 81.2 73.3 73.3 70.3 69.3 73.5
[…]
Qwen2.5-32B-InstructDPO+Judger+CL 91.2 82.6 78.6 72.2 69.0 78.7 91.2 85.1 84.1 79.5 80.4 84.1
Table 4: The overall performance on the soft constraint subset of FollowBench. We use bold for the best results and underlined for the second-best results.
Models Easy Set Hard Set Full Set Avg.
CSR ISR PSR CSR ISR PSR CSR ISR PSR
BASE 0.770 0.450 0.530 0.640 0.160 0.320 0.700 0.300 0.430 0.478
DPO+judger+CL 0.790 0.480 0.560 0.670 0.190 0.320 0.730 0.330 0.440 0.501
Table 5: The overall performance of LLaMA3-8B-Instruct on CFBench. We use bold for the best results.
@@ 4.2 Main Results
Our method (DPO+Judger+CL) significantly enhances the model’s ability to follow soft constraints across different models. Moreover, our method is also effective for hard constraints. As shown in Tab. 4, on FollowBench, which includes both soft and hard constraints, our method improves the model’s performance. The improvement is significant on difficult constraint following tasks, especially at the L5 difficulty level in FollowBench. Specifically, LLaMA-3-8B-Instruct shows an improvement of 5.5% at the L5 difficulty level. For soft constraints, as shown in Tab. 4.1 and Tab. 5, we validate the effectiveness of our method using the soft constraint subset of FollowBench and CFBench. The results show that our method is also effective in improving performance on soft constraints. For hard constraints, model’s performance on IFEval which includes only hard constraints demonstrates the effectiveness of our method. After training with SFT+Judger, there may be a decline in the performance. This drop is attributed to the model’s integration of various specialized training techniques during its initial training phase.
From the perspective of constraint types, as shown in Fig. 4, our method significantly improves the model’s performance across different types of constraints. The most notable improvement is observed in the Mixed category, which is defined as a composition of multiple constraint categories to simulate real-world scenarios. Our method enhances the model’s performance by 10.6% in this category, suggesting a notable enhancement in the model’s ability to handle complex constraints in real-world scenarios. Moreover, our method improves model’s performance by 6% in the Style category which contains only soft constraints. In the category of situation constraints, the decrease in performance is due to the fact that our constructed situation constraints mainly focus on soft constraints. As a result, our method struggles with hard constraints in this category, such as complex situational reasoning.